# Collider imprint of vector-like leptons in light of anomalous magnetic moment and neutrino data

### v1

@@ Abstract
We investigate the impact of incorporating anomaly-free representations of vector-like leptons into the Standard Model, aiming to address longstanding puzzles related to the anomalous magnetic moments of the muon and electron while maintaining consistency with neutrino masses and mixings. We find that among the various representations of vector-like leptons permitted by the Standard Model gauge symmetry, only weak doublets and singlets offer satisfactory solutions, all associated with a significantly constrained parameter space. Our analysis delves into the associated parameter space, identifying representative benchmark scenarios suitable for collider studies. These setups yield a distinctive six-lepton signature whose associated signals can easily be distinguished from the Standard Model background, providing a clear signal indicative of new physics models featuring vector-like leptons. Our work hence sheds light on the potential implications of vector-like leptons in resolving discrepancies inherent to the Standard Model, while also offering insights into experimental avenues for further exploration.
@@ 1 Introduction
$a_{\mu}^{\rm exp}\,(\text{FNAL+BNL})=116~{}592~{}059\,(22)\times 10^{-11}\,(0.19\,\rm ppm)\,.$ (1)
@@
$a_{\mu}^{\rm SM}=116~{}591~{}810\,(43)\times 10^{-11}\,,$ (2)
as evaluated by the Muon $g-2$ Theory Initiative in 2020 Aoyama:2020ynm . However, the recent evaluation of the leading-order hadronic vacuum polarisation contribution to $a_{\mu}$ using lattice simulations suggests a reconciliation between the theoretical and experimental values for $a_{\mu}$ Borsanyi:2020mff ; Ce:2022kxy ; ExtendedTwistedMass:2022jpw ; FermilabLatticeHPQCD:2023jof . These lattice results nevertheless lead to discrepancies with the measurements of properties of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ scattering Wittig:2023pcl , leaving hence the possibility open for new physics contributions. Thus, the $(g-2)_{\mu}$ puzzle continues to motivate the exploration of BSM avenues for restoring agreement between theory and data, as showcased for instance in Jegerlehner:2009ry .
@@
The most recent combined experimental measurements yield Fan:2022eto
@@
$\Delta a_{e}\equiv\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}=\left\{\begin{array}[]{cc}48(30)\times 10^{-14}&{\rm Rb}\\ -88(36)\times 10^{-14}&{\rm Cs}\end{array}\right.~{}.$ (5)
@@
Several BSM extensions propose the existence of new exotic heavy leptons, notably vector-like leptons (VLLs) with identical left-handed and right-handed couplings akin to those originating from composite paradigms Panico:2015jxa ; Cacciapaglia:2020kgq ; Cacciapaglia:2022zwt or present in several neutrino mass models Cai:2017mow . Equipping particle physics theories with VLLs has moreover shown promises in addressing various challenges and anomalies. For instance, BSM setups with VLLs can encompass a dark matter candidate Schwaller:2013hqa ; Bahrami:2016has , and they may contribute to resolving discrepancies in the muon’s anomalous magnetic moment, hence potentially restoring agreement between theory and data Poh:2017tfo ; Chun:2020uzw ; Lee:2022nqz . Additionally, VLLs could predict deviations from lepton universality in $B$-decays Poh:2017tfo , offer an explanation for the Cabibbo angle anomaly Crivellin:2020ebi , and align predictions of the $W$-boson mass with recent experimental measurements from the CDF collaboration Lee:2022nqz .
@@
[…]
@@
In this work, we conduct a comprehensive analysis of models incorporating VLLs irrespective of their representation under the SM gauge symmetry, with a main focus on anomaly-free setups and their consequences at colliders. We explore various BSM constructions with VLL singlets, doublets and triplets, and we probe their potential role in addressing fundamental physics puzzles such as dark matter, the muon and electron anomalous magnetic moments, and neutrino masses (all of these corresponding to limitations of the SM that could be related to the existence of an extended leptonic sector). Despite introducing only one or multiple VLL representations, we find no common explanation that could account for all these phenomena. While the idea of dark matter being associated with a vector-like neutrino has been extensively discussed in prior literature Jegerlehner:2009ry , our emphasis lies on options that provide explanations for discrepancies between theoretical predictions and experimental measurements in the context of the lepton anomalous magnetic moments, and that are suitable with respect to neutrino mass and mixing constraints. Additionally, we ensure that our solutions are compatible with electroweak precision data, and finally we investigate their implications at collider experiments.
@@ 2 Theoretical Framework
Unlike SM-like chiral fermions, whose left-handed and right-handed components transform differently under the SM gauge symmetry $SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}$, vector-like fermions exhibit identical properties for both their chiralities. However, the specific choice of representation for these fermions is not unique. In models featuring VLLs, building a consistent, renormalisable, and anomaly-free extension of the SM imposes constraints on the possible choices for their $SU(2)_{L}$ and hypercharge quantum numbers.
@@
Table 1: Possible VLL representations leading to an anomaly-free and renormalisable BSM theory. We consider weak singlets (${\cal S}_{1}$ and ${\cal S}_{2}$), doublets (${\cal D}_{1}$, ${\cal D}_{2}$), and triplets (${\cal X}_{1}$ and ${\cal X}_{2}$) with different hypercharge quantum number assignments. In the second line of the table, we provide the components of the various weak multiplets considered (with their electric charge explicitly indicated), while the $SU(2)_{L}\times U(1)_{Y}$ quantum numbers are given in the last lines of the table.
@@
To address the anomalous magnetic moment of both the electron and the muon as well as other puzzles of the SM such as neutrino masses and mixings, we construct a theoretical framework that incorporates one or more of the considered VLL representations, each endowed with a generation index. Furthermore, we enforce that a given VLL species only interacts with the corresponding generation of ordinary leptons. This approach, distinct from introducing VLL fields with inter-generational couplings, allows us to analyse the impact of the model on the anomalous magnetic moments of the electron and muon separately, in a way that avoids inducing lepton-violating decays. In our new physics setup, the VLL collider signatures and their corresponding rates (including production cross sections and relevant branching ratios) depend on how the different VLLs mix with ordinary leptons, which stems from their Yukawa interaction with the Higgs field $\Phi$, as well as their mass terms. The corresponding Lagrangian contributions, with generational indices understood, are given by
@@
$\begin{split}&{\cal L}_{\mathrm{Y}}\!=-\!\bigg{[}\mathbf{Y_{e}}\,{\bar{L}}\Phi e+\mathbf{Y_{\nu}}\,({\bar{L}}\cdot\Phi^{\dagger})\nu+\mathbf{\hat{Y}_{\mathcal{S}_{1}}}\,{\bar{L}}\Phi{\cal S}_{1}+\mathbf{\hat{Y}_{\mathcal{S}_{2}}}\,({\bar{L}}\!\cdot\!\Phi^{\dagger}){\cal S}_{2}+\mathbf{\hat{Y}_{\mathcal{D}_{1}}}\,{\bar{\cal D}}_{1}\Phi e+\mathbf{\hat{Y}_{\mathcal{D}_{2}}}\,({\bar{\cal D}}_{2}\!\cdot\!\Phi^{\dagger})e\\ &\quad+\mathbf{\hat{Y}_{\mathcal{X}_{1}}}\,{\bar{L}}{\cal X}_{1}\Phi+\mathbf{\hat{Y}_{\mathcal{X}_{2}}}\,\Phi\!\cdot\!({\cal X}_{2}L)+\mathrm{H.c.}\bigg{]}-\sum_{i=1}^{2}\bigg{[}M_{\mathcal{S}_{i}}\,{\bar{\cal S}}_{i}{\cal S}_{i}+M_{\mathcal{D}_{i}}\,{\bar{\cal D}}_{i}{\cal D}_{i}+M_{\mathcal{X}_{i}}\,{\bar{\cal X}}_{i}{\cal X}_{i}\bigg{]}\,.\end{split}$ (6)
@@
Here, $L$, $e$ and $\nu$ represent the SM doublet and singlet of ordinary leptons and neutrinos [^1] While strictly speaking there is no right-handed neutrino $\nu\equiv\nu_{R}$ in the SM, we consider a SM sector containing such a state, that is thus differently tagged from the VLL representations studied., respectively, and the explicit invariant products appearing in the Lagrangian denote $SU(2)$-invariant products of two fields lying in the fundamental or antifundamental representation of $SU(2)_{L}$. The different parameters appearing in the Lagrangian ${\cal L}_{\mathrm{Y}}$ include the $3\times 3$ SM Yukawa matrices in the flavour space $\mathbf{Y_{e}}$, $\mathbf{Y_{\nu}}$, $\mathbf{\hat{Y}_{\mathcal{S}_{1}}}$, $\mathbf{\hat{Y}_{\mathcal{S}_{2}}}$, $\mathbf{\hat{Y}_{\mathcal{D}_{1}}}$, $\mathbf{\hat{Y}_{\mathcal{D}_{2}}}$, $\mathbf{\hat{Y}_{\mathcal{X}_{1}}}$ and $\mathbf{\hat{Y}_{\mathcal{X}_{2}}}$, along with the mass parameters associated with the VLLs $M_{\mathcal{S}_{1}}$, $M_{\mathcal{S}_{2}}$, $M_{\mathcal{D}_{1}}$, $M_{\mathcal{D}_{2}}$, $M_{\mathcal{X}_{1}}$ and $M_{\mathcal{X}_{2}}$. All these matrices are imposed to be flavour diagonal, which ensures that a vector-like representation of a given generation only interacts with the SM lepton of the same generation.
@@
$\begin{split}{\cal L}_{\mathrm{VLL}}=&\ \bigg{[}\big{(}{\bar{\cal D}}_{1}\Phi\big{)}\big{(}\lambda_{L}^{(1)}P_{L}+\lambda^{(1)}_{R}P_{R}\big{)}{\cal S}_{1}+\big{(}{\bar{\cal D}}_{1}\cdot\Phi^{\dagger}\big{)}\big{(}\lambda_{L}^{(2)}P_{L}+\lambda^{(2)}_{R}P_{R}\big{)}{\cal S}_{2}+\mathrm{H.c.}\bigg{]}\,.\end{split}$ (7)
@@
Observable Constraints Observable Constraints
$m_{h}$ [122, 128] GeV BR$(B^{0}_{s}\rightarrow\mu^{+}\mu^{-})$ $[1.1,6.4]\times 10^{-9}$
BR$(h\rightarrow\mu^{+}\mu^{-})$ (strength) $[0.51,1.87]$ $\frac{\mathrm{BR}(B\rightarrow\tau\nu_{\tau})}{\mathrm{BR}_{{\rm SM}}(B\rightarrow\tau\nu_{\tau})}$ [0.15, 2.41]
BR$(H\rightarrow\gamma\gamma)$ (strength) $[0.96,1.24]$ $\mathrm{BR}(B^{0}\rightarrow\chi_{s}\gamma)$ $[2.99,3.87]\times 10^{-4}$
$m_{\nu_{1}}+m_{\nu_{2}}+m_{\nu_{3}}$ $\leq 0.26$ eV Neutrino masses See section 3.2
@@
Parameter Scanned range Parameter Scanned range
$|\mathbf{\hat{Y}_{\mathcal{D}_{1}}}|$ $[10^{-5},1]$ $|\mathbf{\hat{Y}_{\mathcal{S}_{1}}}|$ $[10^{-5},1]$
$|\mathbf{\hat{Y}_{\mathcal{S}_{2}}}|$ $[10^{-7},10^{-3}]$ $|\mathbf{Y_{\nu}}|$ $[10^{-15},10^{-10}]$
$|\lambda_{L}^{(1)}|\,,\ |\lambda_{R}^{(1)}|$ $[10^{-5},1]$ $|\lambda_{L}^{(2)}|\,,\ |\lambda_{R}^{(2)}|$ $[10^{-7},10^{-3}]$
$M_{\mathcal{S}_{1}}$ [100, 500] GeV $M_{\mathcal{S}_{2}}$ [100, 1000]
@@
$\Delta a^{\text{Cs}}_{e}\,\equiv\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}\,=\,[\,-168,\,-108\,]\times 10^{-14}\,,\qquad\Delta a^{\text{Rb}}_{e}\,\equiv\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}\,=\,[\,2,\,66\,]\times 10^{-14}\quad\,.$ (9)

### v2

@@ Abstract
We investigate the impact of incorporating vector-like leptons into the Standard Model, aiming to address longstanding puzzles related to the anomalous magnetic moments of the muon and electron while maintaining consistency with neutrino masses and mixings. We find that among the various representations of vector-like leptons permitted by the Standard Model gauge symmetry, only weak doublets and singlets offer satisfactory solutions, all associated with a significantly constrained parameter space. Our analysis delves into the associated parameter space, identifying representative benchmark scenarios suitable for collider studies. These setups yield a distinctive six-lepton signature whose associated signals can easily be distinguished from the Standard Model background, providing a clear signal indicative of new physics models featuring vector-like leptons. Our work hence sheds light on the potential implications of vector-like leptons in resolving discrepancies inherent to the Standard Model, while also offering insights into experimental avenues for further exploration.
journal: Eur. Phys. J. C
@@ 1 Introduction
$a_{\mu}^{\rm exp}\,(\text{FNAL+BNL})=116\leavevmode\nobreak\ 592\leavevmode\nobreak\ 059\,(22)\times 10^{-11}\,(0.19\,\rm ppm)\,.$ (1)
@@
[…]
as evaluated by the Muon $g-2$ Theory Initiative in 2020 Aoyama:2020ynm . However, the recent evaluation of the leading-order hadronic vacuum polarisation contribution to $a_{\mu}$ using lattice simulations suggests a reconciliation between the theoretical and experimental values for $a_{\mu}$ Borsanyi:2020mff ; Ce:2022kxy ; ExtendedTwistedMass:2022jpw ; RBC:2023pvn ; FermilabLatticeHPQCD:2023jof . These lattice results nevertheless lead to discrepancies with the measurements of properties of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ scattering Wittig:2023pcl , leaving hence the possibility open for new physics contributions. Thus, the $(g-2)_{\mu}$ puzzle continues to motivate the exploration of BSM avenues for restoring agreement between theory and data, as showcased for instance in Jegerlehner:2009ry .
@@
The most recent combined experimental measurements, determined in Fan:2022eto , yield
@@
$\Delta a_{e}\equiv\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}=\left\{\begin{array}[]{cc}48(30)\times 10^{-14}&{\rm Rb}\\ -88(36)\times 10^{-14}&{\rm Cs}\end{array}\right.\leavevmode\nobreak\ .$ (5)
@@
Several BSM extensions propose the existence of new exotic heavy leptons, notably vector-like leptons (VLLs) with identical left-handed and right-handed couplings akin to those originating from composite paradigms Panico:2015jxa ; Cacciapaglia:2020kgq ; Cacciapaglia:2022zwt or present in several neutrino mass models Cai:2017mow . Equipping particle physics theories with VLLs has moreover shown promises in addressing various challenges and anomalies. For instance, BSM setups with VLLs can encompass a dark matter candidate Schwaller:2013hqa ; Bahrami:2016has , and they may contribute to resolving discrepancies in the muon’s anomalous magnetic moment, hence potentially restoring agreement between theory and data Poh:2017tfo ; Crivellin:2018qmi ; Hiller:2019mou ; Hiller:2020fbu ; Chun:2020uzw ; De:2021crr ; Escribano:2021css ; Lee:2022nqz . Additionally, VLLs could predict deviations from lepton universality in $B$-decays Poh:2017tfo , offer an explanation for the Cabibbo angle anomaly Crivellin:2020ebi , and align predictions of the $W$-boson mass with recent experimental measurements from the CDF collaboration Lee:2022nqz .
@@
One crucial motivation for including VLLs in theoretical frameworks is their potential to extend the SM neutrino sector to accommodate a variety of experimental data. Neutrinos, among the most mysterious particles in the universe, play essential roles in understanding weak interactions, as well as cosmology, nuclear physics and astrophysics. While expected to be massless in the SM, neutrinos have been shown to oscillate between flavours, indicating that they have non-zero masses. While the neutrino absolute masses are not known, neutrino oscillations measured mass-square differences, found to be $m_{\nu_{2}}^{2}-m_{\nu_{1}}^{2}=7.5\times 10^{-5}$ eV2 and $m_{\nu_{3}}^{2}-m_{\nu_{1}}^{2}=\pm 2.5\times 10^{-3}$ eV2 deGouvea:2016qpx . Additionally, cosmic surveys indirectly constrain the sum of neutrino masses to be at the sub-eV level, $\sum_{j}m_{\nu_{j}}<0.23$ eV at the 95% confidence level Loureiro:2018pdz . The discovery of non-zero neutrino masses is probably among the most important particle physics results of the last two decades, as it suggests that the SM is incomplete. Thus any consistent model of BSM physics should be able to provide a mechanism and an explanation for neutrino masses and mixings.
@@
In this work, we conduct a comprehensive analysis of models incorporating VLLs irrespective of their representation under the SM gauge symmetry, with a main focus on their consequences at colliders. We explore various BSM constructions with VLL singlets, doublets and triplets, and we probe their potential role in addressing fundamental physics puzzles such as dark matter, the muon and electron anomalous magnetic moments, and neutrino masses (all of these corresponding to limitations of the SM that could be related to the existence of an extended leptonic sector). Despite introducing only one or multiple VLL representations, we find no common explanation that could account for all these phenomena. While the idea of dark matter being associated with a vector-like neutrino has been extensively discussed in prior literature Jegerlehner:2009ry , our emphasis lies on options that provide explanations for discrepancies between theoretical predictions and experimental measurements in the context of the lepton anomalous magnetic moments, and that are suitable with respect to neutrino mass and mixing constraints. Additionally, we ensure that our solutions are compatible with electroweak precision data, and finally we investigate their implications at collider experiments.
The connection between the anomalous magnetic moment of the muon and electron, along with the presence of vector-like leptons in the theory, has been previously studied in several studies. In Kannike:2011ng , old constraints originating from the puzzle associated with the anomalous magnetic moment of the muon were used to define interesting collider signals. Furthermore, Ref. Arcadi:2021cwg recently analysed the connection between $(g-2)_{\mu}$, dark matter and $B$-meson anomalies independently of collider studies, in complementarity with explorations of the impact of vector-like leptons on $(g-2)_{\mu}$ and its correlations with the properties of the Higgs sector, and either leptonic flavour violation Poh:2017tfo , a dark matter candidate Lu:2021vcp , the measurement of the $W$-boson mass by the CDF experiment CDF:2022hxs ; deGiorgi:2022xhr ; Lee:2022nqz , or even in an effective field theory interpretation DelleRose:2022ygn . However, none of these works deeply investigated the consequences on potential VLL signals at colliders by means of a state-of-the-art simulation tool chain. Non-minimal fermion-scalar TeV-scale extensions of the SM were also explored, but again without examining any VLL collider implication Guedes:2022cfy . In addition, a joint explanation to observed deviations in both the electron and muon anomalous magnetic moments was discussed in Craig:2021ksw through the identification of a hidden symmetry responsible for a “magic zero” implying that the leading contribution to the muon $g-2$ vanishes upon integrating out weak doublet and singlet of vector-like fermions. Non-minimal model studies also include setups in which the SM is augmented not only by vector-like leptons, but also by additional gauge symmetries Kawamura:2022fhm ; Zhou:2022cql or an extended scalar sector Chun:2020uzw ; Raju:2022zlv ; Frank:2020smf ; DeJesus:2020yqx ; Bharadwaj:2021tgp ; Chakrabarty:2020jro , as well as in warped space constructions Megias:2017dzd .
In comparison, our work aims to bring two new important features complementing previous explorations. First, our analysis includes the most up-to-date constraints from both BNL and FNAL experiments for the anomalous magnetic moment of the muon, and second, it includes a comprehensive collider analysis for promising benchmarks which are consistent with both satisfying constraints from low energy data, neutrino masses and mixings, and anomalous moment measurements of both the muon and the electron.
@@ 2 Theoretical Framework
Unlike SM-like chiral fermions, whose left-handed and right-handed components transform differently under the SM gauge symmetry $SU(3)_{C}\otimes SU(2)_{L}\otimes U(1)_{Y}$, vector-like fermions exhibit identical properties for both their chiralities. However, the specific choice of representation for these fermions is not unique. In models featuring VLLs, building a consistent, renormalisable, and anomaly-free (as automatically guaranteed by the addition of vector-like fermions) extension of the SM imposes constraints on the possible choices for their $SU(2)_{L}$ and hypercharge quantum numbers, especially once we consider only gauge-covariant lepton multiplets which can mix with the ordinary leptons through renormalisable couplings. In other words, we will only consider interactions with ordinary fermions through Yukawa couplings to the SM scalar Aguilar-Saavedra:2013qpa.
@@
Table 1: Possible VLL representations leading to a renormalisable BSM theory. We consider weak singlets (${\cal S}_{1}$ and ${\cal S}_{2}$), doublets (${\cal D}_{1}$, ${\cal D}_{2}$), and triplets (${\cal X}_{1}$ and ${\cal X}_{2}$) with different hypercharge quantum number assignments. In the second line of the table, we provide the components of the various weak multiplets considered (with their electric charge explicitly indicated), while the $SU(2)_{L}\times U(1)_{Y}$ quantum numbers are given in the last lines of the table.
@@
To address the anomalous magnetic moment of both the electron and the muon as well as other puzzles of the SM such as neutrino masses and mixings, we construct a theoretical framework that incorporates one or more of the considered VLL representations, each endowed with a generation index. In other words, we enforce that a given VLL species primarily interacts with the corresponding generation of ordinary leptons, and next neglect any inter-generational coupling. This approach, distinct from introducing VLL fields with inter-generational couplings, allows us to analyse the impact of the model on the anomalous magnetic moments of the electron and muon separately, in a way that avoids inducing lepton-flavour-violating (LFV) decays. While it is arguable that inter-generational mixing is generated nevertheless through quantum corrections, LFV constraints enforce it to be small. The VLL collider signatures of our setup and their corresponding rates (including production cross sections and relevant branching ratios) depend on how the different VLLs mix with ordinary leptons, which stems from their Yukawa interaction with the Higgs field $\Phi$, as well as their mass terms. The corresponding Lagrangian contributions, with generational indices understood, are given by
@@
$\begin{split}&{\cal L}_{\mathrm{Y}}=-\bigg{[}\mathbf{Y_{e}}\,{\bar{L}}\Phi e+\mathbf{Y_{\nu}}\,({\bar{L}}\cdot\Phi^{\dagger})\nu+\mathbf{\hat{Y}_{\mathcal{S}_{1}}}\,{\bar{L}}\Phi{\cal S}_{1}\\ &\quad+\mathbf{\hat{Y}_{\mathcal{S}_{2}}}\,({\bar{L}}\!\cdot\!\Phi^{\dagger}){\cal S}_{2}+\mathbf{\hat{Y}_{\mathcal{D}_{1}}}\,{\bar{\cal D}}_{1}\Phi e+\mathbf{\hat{Y}_{\mathcal{D}_{2}}}\,({\bar{\cal D}}_{2}\!\cdot\!\Phi^{\dagger})e\\ &\quad+\mathbf{\hat{Y}_{\mathcal{X}_{1}}}\,{\bar{L}}{\cal X}_{1}\Phi+\mathbf{\hat{Y}_{\mathcal{X}_{2}}}\,\Phi\!\cdot\!({\cal X}_{2}L)+\mathrm{H.c.}\bigg{]}\\ &\quad-\sum_{i=1}^{2}\bigg{[}M_{\mathcal{S}_{i}}\,{\bar{\cal S}}_{i}{\cal S}_{i}+M_{\mathcal{D}_{i}}\,{\bar{\cal D}}_{i}{\cal D}_{i}+M_{\mathcal{X}_{i}}\,{\bar{\cal X}}_{i}{\cal X}_{i}\bigg{]}\,.\end{split}$ (6)
@@
Here, $L$, $e$ and $\nu$ represent the SM doublet and singlet of ordinary leptons and neutrinos [^1] While strictly speaking there is no right-handed neutrino $\nu\equiv\nu_{R}$ in the SM, we consider a SM sector containing such a state, that is thus differently tagged from the VLL representations studied. However, all bilinear terms involving a right-handed neutrino have been ignored for simplicity, as they do not impact our ability to find scenarios compatible with all constraints considered in this work and the associated collider phenomenology., respectively, and the explicit invariant products appearing in the Lagrangian denote $SU(2)$-invariant products of two fields lying in the fundamental or anti-fundamental representation of $SU(2)_{L}$. The different parameters appearing in the Lagrangian ${\cal L}_{\mathrm{Y}}$ include the $3\times 3$ SM Yukawa matrices in the flavour space $\mathbf{Y_{e}}$, $\mathbf{Y_{\nu}}$, $\mathbf{\hat{Y}_{\mathcal{S}_{1}}}$, $\mathbf{\hat{Y}_{\mathcal{S}_{2}}}$, $\mathbf{\hat{Y}_{\mathcal{D}_{1}}}$, $\mathbf{\hat{Y}_{\mathcal{D}_{2}}}$, $\mathbf{\hat{Y}_{\mathcal{X}_{1}}}$ and $\mathbf{\hat{Y}_{\mathcal{X}_{2}}}$, along with the mass parameters associated with the VLLs $M_{\mathcal{S}_{1}}$, $M_{\mathcal{S}_{2}}$, $M_{\mathcal{D}_{1}}$, $M_{\mathcal{D}_{2}}$, $M_{\mathcal{X}_{1}}$ and $M_{\mathcal{X}_{2}}$. All these matrices are imposed to be flavour diagonal, which ensures that a vector-like representation of a given generation only interacts with the SM lepton of the same generation. Furthermore, terms involving different VLL species have all been omitted for now, and will be discussed below.
@@
$\begin{split}{\cal L}_{\mathrm{VLL}}=&\ \bigg{[}\big{(}{\bar{\cal D}}_{1}\Phi\big{)}\big{(}\lambda_{L}^{(1)}P_{L}+\lambda^{(1)}_{R}P_{R}\big{)}{\cal S}_{1}\\ &\quad+\big{(}{\bar{\cal D}}_{1}\cdot\Phi^{\dagger}\big{)}\big{(}\lambda_{L}^{(2)}P_{L}+\lambda^{(2)}_{R}P_{R}\big{)}{\cal S}_{2}+\mathrm{H.c.}\bigg{]}\,.\end{split}$ (7)
@@
Observable Constraints
$m_{h}$ [122, 128] GeV
BR$(h\rightarrow\mu^{+}\mu^{-})$ (strength) $[0.51,1.87]$
BR$(h\rightarrow\gamma\gamma)$ (strength) $[0.96,1.24]$
BR$(B^{0}_{s}\rightarrow\mu^{+}\mu^{-})$ $[1.1,6.4]\times 10^{-9}$
$\mathrm{BR}(B^{0}\rightarrow\chi_{s}\gamma)$ $[2.99,3.87]\times 10^{-4}$
$\frac{\mathrm{BR}(B\rightarrow\tau\nu_{\tau})}{\mathrm{BR}_{{\rm SM}}(B\rightarrow\tau\nu_{\tau})}$ [0.15, 2.41]
$m_{\nu_{1}}+m_{\nu_{2}}+m_{\nu_{3}}$ $\leq 0.26$ eV
Neutrino masses See section 3.2
@@
      Parameter       Scanned range
      $|\mathbf{\hat{Y}_{\mathcal{D}_{1}}}|$       $[10^{-5},1]$
      $|\mathbf{\hat{Y}_{\mathcal{S}_{1}}}|$       $[10^{-5},1]$
      $|\mathbf{\hat{Y}_{\mathcal{S}_{2}}}|$       $[10^{-7},10^{-3}]$
      $|\mathbf{Y_{\nu}}|$       $[10^{-15},10^{-10}]$
      $|\lambda_{L}^{(1)}|\,,\ |\lambda_{R}^{(1)}|$       $[10^{-5},1]$
      $|\lambda_{L}^{(2)}|\,,\ |\lambda_{R}^{(2)}|$       $[10^{-7},10^{-3}]$
      $M_{\mathcal{S}_{1}}$       [100, 500] GeV
      $M_{\mathcal{S}_{2}}$       [100, 1000]
@@
Furthermore, it is important to note that VLL mixings with the SM lepton fields entail modifications of the couplings of the $W$ and $Z$ bosons with the SM-like leptons $\ell$. The relevant neutral-current Lagrangian terms can be expressed as follows, considering two SM leptons with flavour indices $i$ and $j$ (where $i,j=1,2,3$ correspond to the electron, muon and tau, respectively):
[…]
where $L_{\ell}=g/c_{W}(-1/2+s_{W}^{2})$ and $R_{\ell}=g/c_{W}s_{W}^{2}$ represent the SM $Z$-boson couplings to left-handed and right-handed leptons. Here, $s_{W}$ and $c_{W}$ denote the sine and cosine of the electroweak mixing angle, respectively, and $g$ stands for the weak coupling constant. This interaction Lagrangian explicitly depends on the lepton mixing matrices $U^{L}_{\ell}$ and $U^{R}_{\ell}$ that diagonalise the mass matrix
$m_{\ell}=\begin{pmatrix}\frac{v}{\sqrt{2}}\mathbf{Y_{e}}^{\!\!T}&0&\frac{v}{\sqrt{2}}\mathbf{\hat{Y}_{\mathcal{S}_{1}}}\\ \frac{v}{\sqrt{2}}\mathbf{\hat{Y}_{\mathcal{D}_{1}}}&M_{\mathcal{D}_{1}}&\frac{v}{\sqrt{2}}\lambda_{R}^{(1)}\\ 0&\frac{v}{\sqrt{2}}\lambda_{L}^{(1)\,T}&M_{\mathcal{S}_{1}}\end{pmatrix}\,,$ (9)
such that
${\rm diag}(m_{e},m_{\mu},\ldots)=U^{L\,*}_{\ell}\,m_{\ell}\,U^{R\,\dagger}_{\ell}\,.$ (10)
In this expression, $v$ represents the vacuum expectation value of the SM Higgs doublet. Due to the significant differences in the orders of magnitude among the parameters present in the mass matrix, inter-generational mixing for the ordinary leptons is found to be negligible. Consequently, the couplings in Eq. (8) that involve ordinary leptons can be approximated to those of the SM. For example, in the case of the $Z$-boson coupling to the electron, we obtain:
[…]
Thus, the model examined in this study aligns with the established properties of the $Z$ boson. Similarly, the charged-current Lagrangian involving $W$-boson couplings to an ordinary charged lepton $\ell_{j}$ and neutrino $\nu_{i}$ (for $i=1,2,3$ being a flavour index) is given by
[…]
This Lagrangian depends both on the matrices $U^{L}_{\ell}$ and $U^{R}_{\ell}$ diagonalising the charged lepton sector, and on the matrices $U_{\nu}^{L}$ and $U_{\nu}^{R}$ that diagonalise the neutrino mass matrix,
$m_{\nu}=\begin{pmatrix}\frac{1}{\sqrt{2}}v\mathbf{Y_{\nu}}&0&\frac{1}{\sqrt{2}}v\mathbf{\hat{Y}_{\mathcal{S}_{2}}}\\ 0&M_{\mathcal{D}_{1}}&\frac{1}{\sqrt{2}}v\lambda_{R}^{(2)}\\ 0&0&M_{\mathcal{S}_{2}}\end{pmatrix}$ (13)
such that
${\rm diag}(m_{\nu_{1}},m_{\nu_{2}},\ldots)=U_{\nu}^{L\,*}m_{\nu}U_{\nu}^{R\,\dagger}\,.$ (14)
Once again, by virtue of the typical values of the different parameters entering that mass matrix, the $W$-boson couplings to SM leptons and neutrinos can be approximated as in the SM,
[…]
such that the model examined in this work does not challenge measured $W$-boson couplings.
@@
$\begin{split}\Delta a^{\text{Cs}}_{e}\,\equiv&\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}\,=\,[\,-168,\,-108\,]\times 10^{-14}\,,\\ \Delta a^{\text{Rb}}_{e}\,\equiv&\,a^{\text{exp}}_{e}-a^{\text{SM}}_{e}\,=\,[\,2,\,66\,]\times 10^{-14}\quad\,.\end{split}$ (17)
There is some controversy in the literature regarding whether the anomalous magnetic moment of the muon constitutes a signal of physics beyond the SM. Recent lattice simulations of the hadronic vacuum polarisation contribution to $a_{\mu}$ have suggested a reconciliation between theoretical predictions and experimental measurements Borsanyi:2020mff ; Ce:2022kxy ; ExtendedTwistedMass:2022jpw ; RBC:2023pvn ; FermilabLatticeHPQCD:2023jof , while also revealing discrepancies in the properties of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ scattering Wittig:2023pcl . As a result, while the likelihood that there are large new physics contributions to $a_{\mu}$ has diminished, some unresolved issues remain and continue to motivate BSM explorations in this area. This leaves open the question of whether the deviation in Eq. (16) would still be considered a sign of new physics affecting hadronic vacuum polarisation. If this interpretation holds, it is important to remember that such contributions generally worsen electroweak precision fits. Conversely, models which predict smaller contributions to $a_{\mu}$ may become increasingly attractive in light of the recent lattice findings (and a negative answer to the above question). For an up-to-date discussion, see Aoyama:2020ynm .